OpenStudy (anonymous):

What is the area of a circle with a diameter of 8? (1 point) 4 pi 8 pi 16 pi 64 pi

OpenStudy (anonymous):

I need help with this one and 3 more, if anyone would like to help.

OpenStudy (anonymous):

area of a circle is \(\pi r^2\) if your diameter is 8, then the radius is 4, so the area is \(\pi \times 4^2=16\pi\)